Three numbers are in the ratio 2:4:7. If the first number is multiplied by 4, the result is 9 less than the sum of the other two. find all three numbers.
Answer by ankor@dixie-net.com(22740)

Three numbers are in the ratio 2:4:7. If the first number is multiplied by 4, the result is 9 less than the sum of the other two. find all three numbers.
:
WE can call it 2x; 4x; 7x without changing the ratio
1st number = 2x
2nd number = 4x
3rd number = 7x
:
Write an equation for:
"If the first number is multiplied by 4, the result is 9 less than the sum of the other two."
4(2x) = 4x + 7x - 9
8x = 11x - 9
8x - 11x = -9
-3x = -9
x = + 3
:
Our three numbers:
2(3) = 6
4(3) = 12
7(3) = 21